OpenStudy (anonymous):

Chem 2 help! 1) write Kc and Kp form if applicable 2) write it backwards. (I guess she means the reverse form) a)HF(aq) + H2O(l) <--> H3O+(aq) + F-(aq) b) 2NO(g) + O2(g) <--> 2NO2(g) c)) CH3COOH(aq) + C2H5OH(aq ) <--> CH3COOC2H5(aq) + H2O(l) I understand how to get Kc and Kp if applicable. But I am not understanding how I would "write it backwards". The teacher told us something about the backwards form being: K (-1)=(1/k) so would that mean K(-1)=(1/Kc) and (1/Kp)?? o.O Please can someone shed some light? :O Please and thank you!

OpenStudy (chmvijay):

see u should find Kc for the non gaseous reaction and Kp for the gaseous reactions

OpenStudy (chmvijay):

KC is = Products / reactants this is for forward reaction if you want to write ir for backward reaction then it becomes KC' =Reactants/ products =1/Kc

OpenStudy (chmvijay):

for example )HF(aq) + H2O(l) <--> H3O+(aq) + F-(aq) Kc= [H3O] [F-] / [H2O] [HF] this is equilibrium constant for forward reaction now let us right for the backward reaction Kc' =[H2O] [HF] / [H3O] [F-] =1/ [H3O] [F-] / [H2O] [HF] =1/Kc Kc = equllibrium constant for forward reaction Kc' =equilibrium constant for forward reaction similarly you can do for other reaction

OpenStudy (anonymous):

Ok, that makes sense! So this only applies to Kc?

OpenStudy (chmvijay):

no you can do for Kp tooo Kp' =1/Kp

OpenStudy (chmvijay):

Kc' =equilibrium constant for forward reaction in the above answer please change forward reaction to backward reaction by mistake i wrote it as forward reaction

OpenStudy (chmvijay):

Kc' =equilibrium constant for backward reaction it should be!!!:)